dontsaymyname:

Can some1 help me pleaseeeeeeeee Its Math

dontsaymyname:

1 attachment
IMTHTBOYTAE:

What is it

ILoveYouLola12:

what is it

dontsaymyname:

Its in the attachment :/

BestBoiKirby:

im dum so probs not

XioGonz:

I can't see the attachment, can you type it out?

dontsaymyname:

I'm dm u it cuz its a lot

XioGonz:

For some reason, I can't see photos in general

darkknight:

hold on, ill get to you in an hour approx. @darkknight

dontsaymyname:

Ok, thanks

darkknight:

also @ramen if/when you are free

jbpeppers809:

2+2 is 4-1 thats 3 quickmath

dontsaymyname:

huh @jbpeppers809

dontsaymyname:

@darkknight Hey jw if ur able to help? I have more questions, but I cant post them with this open.

darkknight:

si, here now

darkknight:

Are you allowed to use a graphing calculator? jw

dontsaymyname:

Yes

dontsaymyname:

I'm going to close this bc i need some help with other questions, is that okay?

darkknight:

alright, u need hlep on this one still or?

dontsaymyname:

Yes I do, please and thank you

darkknight:

Okay, so we have only need 3 points to determine the equation of the quadratic. Lets pick the points 30,29) 40,22) and (50,17) Now we need to know this formula \[ax^2+bx+c = y\] We can plug in the above x and y-coordinates into this equation and solve for a, b and c

darkknight:

So our 3 equations will be \[a(30)^2+b(30)+c = 29\] \[a(40)^2+b(40)+c = 22\] \[a(50)^2+b(50)+c = 17\]

darkknight:

Now we can do some simplifying \[900a+30b+c =29\] \[1600a+40b+c =22\] \[2500a+50b+c =17\]

darkknight:

Now we have to solve for the variables, lets start with eliminating c, so 900a+30b+c=29 1600a+40b+c=22 Applying subtraction -700a -10b = 7 Now we can rearrange ^

darkknight:

Now lets take these 2 eqns 1600a+40b+c=22 2500a+50b+c=17 And eliminate c So we get -900a-10b = 5 so -900a = 5 + 10b or a = (5+10b)/-900

darkknight:

@dontsaymyname

dontsaymyname:

I'll try, thank you soooooooooooooooooo very much @darkknight

darkknight:

We had this eqn previously -700a -10b = 7 Plug in (5+10b)/-900 for a And then \[−700((5+10b)/−900)−10b=7\] simplifying \[((7(5+10b))/9)−10b=7\] so \[(35/9+70b/9)−10b=7\] Can you solve for B?

darkknight:

Anyways, once you get b, plug into a = (5+10b)/-900 OR -700a -10b = 7 You should get the same a value if u try on both. Then plug in both a and b and solve for c

dontsaymyname:

Okayyyy I'll try that, thanks again